Two charge are placed at distance 5km to each other if first charge is displaced 2km toward other charge then after how much time other charge will feel change in force due to first.


1
Expert's answer
2021-02-04T19:09:18-0500

The force change need to travel "5km-2km = 3km = 3\times 10^3m" from the first charge to the second. Since it travels with the speed of light "c = 3\times 10^8m/s" , the time it takes is:


"t = \dfrac{3\times 10^3m}{3\times 10^8m/s} = 10^{-5}s"

Answer. "10^{-5}s."